OpenStudy (anonymous):

Write an equation for the parabola whose vertex is at (2,6)and which passes through(4,-4)

OpenStudy (anonymous):

what is the general form of a parabola?

OpenStudy (anonymous):

Oh crap. I closed my e2020 question down, so I have a different question. Write an equation for the parabola with the given vertex that passes through the given point. Vertex: (-4/3) Point: (-3,6)

OpenStudy (anonymous):

what is the general form of a parabola?

OpenStudy (anonymous):

I have no clue lol... but the choices are a.y=9(x-6)^2+1 b.y=3(x+4)^2+3 c.y=-3(x-5)^2+4 d.y=1/3x^2+5

OpenStudy (anonymous):

then I'm sorry but I can't help you lol.

OpenStudy (anonymous):

aww okay :( thanks for trying lol